OpenStudy (anonymous):

cos(2x) cos(x) - sin(2x) sin(x)...what are the possible solutions for this x is greater than or equal to 0 but less than 360

OpenStudy (cwrw238):

use the identity cos(A +B) = cosAcosB-sinAsinB

OpenStudy (xapproachesinfinity):

i don't see an equation here?

OpenStudy (cwrw238):

no - just noticed that! where's the '=' ??

OpenStudy (xapproachesinfinity):

Shoot! op throw me off the the thread hehe so yes we need = something equal to something for us to solve

OpenStudy (xapproachesinfinity):

\(cos2xcosx-sin2xsinx=cos(3x)\) that's for thr left hand side we need to know the right hand side of the equation? you didn't write it

OpenStudy (cwrw238):

we can't help you any further unless you can give us that

OpenStudy (xapproachesinfinity):

You haven't post the question yet lol

OpenStudy (anonymous):

oh that would be = to 0..sorry I forgot

OpenStudy (xapproachesinfinity):

oh hehehe the equation has to always have to equal sides otherwise it is not an equation lol anyways so we have \(\huge cos2xcosx-sin2xsinx=0\)

OpenStudy (xapproachesinfinity):

As I said above the left hand side is \(\large \rm cos2xcosx-sin2xsinx=cos(2x+x)=cos(3x)\)

OpenStudy (xapproachesinfinity):

I used the identity of sum of two angles like @cwrw238 mentioned above

OpenStudy (xapproachesinfinity):

so i got \(\huge \rm cos(3x)=0\)

OpenStudy (anonymous):

thank you guys, I already got the right answer...

OpenStudy (xapproachesinfinity):

Now solve for x

OpenStudy (xapproachesinfinity):

oh ok! np

OpenStudy (xapproachesinfinity):

Hit best answer thingy if you are satisfied -_^